At 9:00 on Saturday morning, two bicyclists heading in opposite directions pass each other on a bicycle path. The bicyclist heading north is riding 6 km/hour faster than the bicyclist heading south. At 10:15, they are 42.5 km apart. Find the two bicyclists rates.

Answer by lwsshak3(11628)

At 9:00 on Saturday morning, two bicyclists heading in opposite directions pass each other on a bicycle path. The bicyclist heading north is riding 6 km/hour faster than the bicyclist heading south. At 10:15, they are 42.5 km apart. Find the two bicyclists rates.
***
let x=rate of bicyclist heading south
x+6=rate of bicyclist heading north
distance=travel time*rate
..
travel time=1 hr and 15 min=1.25 hr
1 .25x+1.25(x+6)=42.5
1.25x+1.25x+7.5=42.5
2.5x=35
x=14
x+6=20
rate of bicyclist heading south=14 km/hr
rate of bicyclist heading north=20 km/hr